The RBCAB Corporation just paid a dividend of $1.13 per share. The company's CFO expects that the dividend will remain at that level for three years. After year three, it is expected that the dividend will grow at a rate of 4% indefinitely. If the required return is 10%, what is the value of stock today? At 5 years? At 10 years?
Answer #1

a). Value of Stock today = [D1 / (1 + r)] + [D2 / (1 + r)2] + [D3 / (1 + r)3] + [{D3 * (1 + gC)} / {(r - gC) * (1 + r)3}]

= [$1.13 / (1 + 0.10)] + [$1.13 / (1 + 0.10)2] + [$1.13 / (1 + 0.10)3] + [{$1.13 * (1 + 0.04)} / {(0.10 - 0.04) * (1 + 0.10)3}]

= $1.03 + $0.93 + $0.85 + $14.72 = $17.53

b). V5 = D6 / [r - g] = [D3 * (1 + g)3] / [r - g] = [$1.13 * (1 + 0.04)3] / [0.10 - 0.04] = $1.27 / 0.06 = $21.18

c). V10 = V5 * (1 + g)5 = $21.18 * (1 + 0.04)5 = $25.77